# Observable CP-violation in charmed baryons decays with SU(3) symmetry analysis

### v1

@@ Abstract
Introduction- CP violation (CPV) is one of the most important issues in particle physics Lee:1957qs ; Kobayashi:1973fv ; Deshpande:1994ii ; LHCb:2019hro ; Lenz:2020awd ; Schacht:2021jaz ; Bediaga:2022sxw ; Maccolini:2023luv ; He:2024pxh ; Schacht:2024wgc , necessary to explain the baryon-antibaryon asymmetry of the universe. However, CP violation is only observed in meson decays, and no conclusive signal of CP violation in baryon decays has been observed in particle physics experiments. The current observations of CP violation are not sufficient to explain the matter asymmetry of the universe, making the search for new sources of CP violation important.
@@
$\overline{a}_{6}=\frac{1}{2}(-b_{6}-c_{6}-e_{15}+d_{15})+\frac{1}{2}(a+a^{\prime}),$
@@
Since all the diagrams induced by $O_{1,2}$ are included in TDA diagrams $T_{1-7}$, it is reasonable to believe the diagrams $T_{8,9}$ are induced by other operators like the penguin operators $O_{3-6}$. The inclusion of these additional diagrams suggests the possibility of the involvement of a new weak phase in these processes. In Eq. Observable CP-violation in charmed baryons decays with SU(3) symmetry analysis, we find that $g^{c}_{15}=0$, which is consistent with our global fit (Case I) in Table. 1. One can also directly observe that $f^{b}_{6}$ depends only on new effect form factors, which provides us an ideal condition to study whether a different weak phase exists in these processes.

### v2

@@ Abstract
Introduction- CP violation (CPV) is one of the most important issues in particle physics Lee:1957qs ; Kobayashi:1973fv ; Deshpande:1994ii ; LHCb:2019hro ; Lenz:2020awd ; Schacht:2021jaz ; Bediaga:2022sxw ; Wang:2022tcm ; Wang:2022fih ; Maccolini:2023luv ; Shen:2023eln ; Song:2024jjn ; He:2024pxh ; Schacht:2024wgc ; Wang:2024ztg , necessary to explain the baryon-antibaryon asymmetry of the universe. However, CP violation is only observed in meson decays, and no conclusive signal of CP violation in baryon decays has been observed in particle physics experiments. The current observations of CP violation are not sufficient to explain the matter asymmetry of the universe, making the search for new sources of CP violation important.
@@
$\overline{a}_{6}=\frac{1}{2}(-b_{6}-c_{6}-e_{15}+d_{15})+\frac{1}{4}(a+a^{\prime}),$
@@
Since all the diagrams induced by $O_{1,2}$ are included in TDA diagrams $T_{1-7}$, it is reasonable to believe the diagrams $T_{8,9}$ are induced by SU(3) symmetry breaking or other operators like the penguin operators $O_{3-6}$. The inclusion of these additional diagrams suggests the possibility of the involvement of a new weak phase in these processes. In Eq. Observable CP-violation in charmed baryons decays with SU(3) symmetry analysis, we find that $g^{c}_{15}=0$, which is consistent with our global fit (Case I) in Table. 1. One can also directly observe that $f^{b}_{6}$ depends only on new effect form factors, which provides us an ideal condition to study whether a different weak phase exists in these processes.